# A uniform-in-time nonlocal approximation of the standard Fokker-Planck equation

### v2

@@ 3.1 Harris’s Theorem
where $L^{*}$ is the (formal) dual operator of $L$ acting on “nice enough” functions. The latter condition has the technical burden of appropriately defining $L^{*}$ (see for example Hairer (2010)), where inequality (14) must be understood in a submartingale sense). Our approach is really the same as this, but we prefer to use Hypothesis 1, which does not have any technical issues. We actually show condition (14), in a suitable sense, for the generator associated to equation (NLFP) in Theorems 3.2 and 3.4. From this we will deduce Hypothesis 1.
@@
A stochastic operator $S$ on $\mathcal{M}$ satisfies a Harris condition on a set $\mathcal{S}\subseteq\Omega$ if there exist $10<\alpha<$ 1 and $\rho\in\mathcal{P}$ such that
@@
$\left\|S_{t}\nu\right\|_{V}\leq Ce^{-\lambda t}\left\|\nu\right\|_{V}\quad\text{ for all $\nu\in\mathcal{P}_{V}$ with $\int_{\Omega}\nu=0$ and all $t\geq 0$},$
@@ 3.2 Lyapunov condition
$L_{\varepsilon}^{*}\varphi:=A^{*}_{\varepsilon}\varphi+B_{\varepsilon}^{*}\varphi=\frac{1}{\varepsilon^{2}}[J_{\varepsilon}*\varphi-\varphi]-x\nabla\varphi.$
@@
To lighten the notation in the following, we will use $C_{1},C_{2}$ as general constants, whose value may change from line to line. Recall from (1), that for every $\alpha$ such that $|\alpha|=1$,
$\int J(z)z^{\alpha}\,\mathrm{d}z=0$
@@
where $R_{\alpha}(\xi)$ is the Lagrange reminder for some $\xi(x,z)\in$ $[x,x+\varepsilon z]$, the segment that joins $x$ and $x+\varepsilon z$.
@@
$\partial_{i}^{2}\left\langle x\right\rangle^{k}=k\big{(}\left\langle x\right\rangle^{k-2}+(k-2)x_{i}^{2}\left\langle x\right\rangle^{k-4}\big{)}$ (15)
For $k\in(0,2)$ we have $\partial_{i}^{2}\varphi(\xi)\leq k$, for every $i,\xi$ Hence
$\frac{1}{\varepsilon^{2}}\int_{\mathbb{R}^{d}}J_{\varepsilon}(y)(\left\langle x+y\right\rangle^{k}-\left\langle x\right\rangle^{k})\,\mathrm{d}y\leq\frac{k}{2}\int_{\mathbb{R}^{d}}J(z)|z|^{2}\,\mathrm{d}z=k$
@@
$\begin{split}\partial_{i}^{2}\varphi(\xi)&\leq k(k-1)\langle\xi\rangle^{k-2}\leq k(k-1)\left\langle|x|+\varepsilon|z|\right\rangle_{k-2}\leq C_{1}(\left\langle x\right\rangle^{k-2}+\left\langle z\right\rangle^{k-2})\end{split}$
@@
$\begin{split}\frac{1}{\varepsilon^{2}}&\int_{\mathbb{R}^{d}}J_{\varepsilon}(y)(\left\langle x+y\right\rangle^{k}-\left\langle x\right\rangle^{k})\,\mathrm{d}y\leq\frac{1}{2}\int_{\mathbb{R}^{d}}J(z)\sum_{i=1}^{d}z_{i}^{2}C_{1}(\langle x\rangle^{k-2}+\langle z\rangle^{k-2})\,\mathrm{d}z\\ &\leq C_{1}\langle x\rangle^{k-2}\frac{1}{2}\int_{\mathbb{R}^{d}}J(z)|z|^{2}\,\mathrm{d}z+C_{1}\int_{\mathbb{R}^{d}}J(z)\langle z\rangle^{k}\,\mathrm{d}z\leq C_{2}\langle x\rangle^{k-2}\end{split}$
@@
We point out that as $k$ approaches 00, the right hand side of (16) goes to 00 as well, since we don’t have Lyapunov confinement in $L^{1}$ without weight.
@@
To lighten the notation in the following, we will use $C_{1},C_{2}$ as general constants, whose value may change from line to line. Again we notice that for every $\alpha$ such that $|\alpha|=1$
$\int J(z)z^{\alpha}=0$
As before, for every $\varphi\in\mathcal{C}^{\infty}$, we Taylor expand around $x$, obtaning
@@
where the Lagrange reminder is evaluated at some $\xi(x,z)\in$ $[x,x-\varepsilon z]$.
@@
$\partial_{i}^{2}\varphi(\xi)=\frac{ae^{a\left\langle\xi\right\rangle}}{\left\langle\xi\right\rangle_{2}}\Big{(}a\xi_{i}^{2}+\left\langle\xi\right\rangle-\frac{\xi_{i}^{2}}{\left\langle\xi\right\rangle}\Big{)}\leq C_{1}e^{a\left\langle\xi\right\rangle}\leq C_{1}e^{a(\left\langle x\right\rangle+\varepsilon\left\langle z\right\rangle)}$
@@
We thus have proved a Lyapunov condition on the generator for $V=\left\langle\cdot\right\rangle^{k}$ and $V=e^{a\left\langle\cdot\right\rangle}$. For these weight functions (and in generally for $\mathcal{C}^{2}$ functions), we see that actually Assumption (14) implies Assumption (SLC), which is the statement of the next Theorem.
@@
Let $u_{0}$ be in $D(L)$ by standard semigroup theory $u(t)\in\mathcal{C}^{1}([0,\infty);L^{1}_{k})$, and consequently we compute
$\begin{split}\frac{\mathrm{d}}{\mathrm{d}t}\int_{\mathbb{R}^{d}}u(t,\cdot)\varphi(x)\,\mathrm{d}x&=\int_{\mathbb{R}^{d}}\frac{\mathrm{d}}{\mathrm{d}t}u(t,\cdot)\varphi(x)\,\mathrm{d}x\\ =&\int_{\mathbb{R}^{d}}(Lu)\varphi(x)\stackrel{{\scriptstyle\mathclap{\mbox{*}}}}{{=}}\int_{\mathbb{R}^{d}}u(L^{*}\varphi)\leq\int_{\mathbb{R}^{d}}u(-\lambda\varphi+C)\end{split}$
where to prove $*\stackrel{{\scriptstyle\mathclap{\mbox{*}}}}{{=}}$ we use classic cutoff arguments with bump functions.
@@
Since the $D(L)$ is dense in $L^{1}_{k}$, the claim is true also for any $u_{0}\in L^{1}_{k}$ by classic density argument. ∎
@@ 3.3 An $L^{\infty}$ version of Berry-Esseen’s central limit theorem
Berry-Esseen-type theorems are quantitative versions of the central limit theorem which provide an estimate on the error between the normal distribution and the scaled law. A result of this kind will be crucial for our proof in Section 3.4 of positivity for the nonlocal Fokker-Planck equation. The first version of this theorem was proved independently by Berry (1941) and Esseen (1942) for real, centred i.i.d. random variables $X_{1},\dots X_{n}$ with ${\textrm{Var}}\,(X_{1})=\sigma^{2}$ and $\mathbb{E}(|X_{1}|^{3})=\rho$, obtaining a result of the type
$|F_{n}(x)-\Phi(x)|\leq\frac{C\rho}{\sqrt{n}\sigma^{3}}$
where $\Phi$ is the cumulative distribution function (c.d.f.) of the normal distribution, $F_{n}$ is the c.d.f. of $\bar{S}_{n}=\frac{1}{\sigma\sqrt{n}}\sum_{i=1}^{n}X_{i}$, and $C$ is a universal constant depending on the dimension $d$. Several generalisations have been developed, proving the same kind of results for not i.i.d. random vectors in $\mathbb{R}^{d}$, as well as improving the precision of the value of $C$. Recent versions can be found in Bentkus (2005) and Raič (2019).
@@
Then there exist an integer $N=N(p)$ and a constant $C_{BE}$ depending on $\frac{L}{l},p,d,\rho_{2+s},\left\|f\right\|_{L^{p}}$ such that $\mathfrak{f}_{n}\in L^{\infty}$ for all $n\geq N$ and
@@
where $G$ is the density of the standard normalised Gaussian.
In order to prove this we first present some preliminary results. The first one is a generalisation of the identity $a^{n}-b^{n}=\sum_{i=1}^{n}(a-b)a^{i-1}b^{n-i}$, that can be easily proven by induction:
@@
$\max_{|\alpha|=2}|D^{\alpha}\hat{f}(x)-D^{\alpha}\hat{f}(y)|\leq L|x+y|^{s},$
with $L:=4\rho_{2+s}$.
@@
The $\mathcal{C}^{2}-$regularity is immediate from Fourier transform properties. We are left to show the Holder-continuity of the derivatives of order $2$, namely, for every multi-index $\alpha$ such that $|\alpha|=2$,
$\begin{split}&\left|D^{\alpha}\hat{f}(\xi)-D^{\alpha}\hat{f}(\eta)\right|=\left|-\widehat{x^{\alpha}f}(\xi)+\widehat{x^{\alpha}f}(\eta)\right|=\left|\int_{\mathbb{R}^{d}}\Big{(}e^{-ix\eta}-e^{-ix\xi})x^{\alpha}f(x)\,\mathrm{d}x\right|\\ &\leq\int_{|x|\leq R}|e^{-ix\xi}-e^{-ix\eta}||x|^{2}|f(x)\,\mathrm{d}x+\int_{|x|>R}|e^{-ix\xi}-e^{-ix\eta}|x|^{2}|f(x)\,\mathrm{d}x\\ &\leq 2|\xi-\eta|\int_{|x|\leq R}|x|^{2-s}|x|^{1-s}f(x)\,\mathrm{d}x+2R^{-s}\int_{|x|>R}|x|^{2+s}f(x)\,\mathrm{d}x\\ &\leq 2\rho_{2+s}(|\xi-\eta|R^{1-s}+R^{-s}).\end{split}$
@@
The last ingredient we need is the following technical Lemma ( that is basically (Hauray and Mischler, 2014, lemma 4.8) with a weaker assumption on moments.
@@
    $\forall\xi\in B_{\delta}(0)\quad|\hat{f}(\xi)|\leq e^{-\frac{|\xi|^{2}}{4}}$
@@
    Assume additionally that $f\in L^{p}$, with $p\in(1,\infty]$. For any $\delta>0$ there exists $\kappa=\kappa(\rho_{2+s},\delta,\left\|f\right\|_{L^{p}})\in(0,1)$ such that
    $\sup_{\left|\xi\right|\geq\delta}|\hat{f}(\xi)|\leq\kappa(\delta).$
@@
    $\begin{split}&\left|\int_{0}^{|\xi|}\Big{(}\frac{\xi}{|\xi|}\Big{)}^{T}\Big{(}H_{\hat{f}}\Big{(}z\frac{\xi}{|\xi|}\Big{)}-H_{\hat{f}}(0)\Big{)}\Big{(}\frac{\xi}{|\xi|}\Big{)}(z-|\xi|)\,\mathrm{d}z\right|\\ &\leq\int_{0}^{|\xi|}\left|\!\left|\!\left|(H_{\hat{f}}\Big{(}z\frac{\xi}{|\xi|}\Big{)}-H_{\hat{f}}(0)\right|\!\right|\!\right||z-|\xi||\,\mathrm{d}z\\ &\leq\int_{0}^{|\xi|}L|z|^{s}|z-|\xi||\,\mathrm{d}z=\frac{L|\xi|^{2+s}}{(s+1)(s+2)},\end{split}$
[…]
    Thus, for small $|\xi|<\delta$ which explicitly depends only on $s$ and $\rho_{2+s}$ we have
    $\left|\hat{f}(\xi)\right|\leq\left|1-\frac{3}{8}|\xi|^{2}\right|\leq e^{-\frac{|\xi|^{2}}{4}}.$
@@
We can now present the proof of our version of Berry-Esseen theorem, that follows the ideas presented in (Goudon et al., 2002, Thm. 1) and (Hauray and Mischler, 2014, Thm 4.6).
@@
First, we notice that $\mathfrak{f}_{n}$ satisfies the hypotheses of Lemma 3.9 for every $n$, and, in particular
$\int_{\mathbb{R}^{d}}\mathfrak{f}_{n}(x)x_{i}x_{k}=\delta_{ij}.$
@@
$\begin{split}{\hat{\mathfrak{f}}_{n}}(\xi)&=\int_{\mathbb{R}^{d}}e^{-ix\xi}(\bar{\sigma}^{2}n)^{d/2}f^{\sigma_{1}\dots,\sigma_{n}}(\sqrt{n}\bar{\sigma}x)\,\mathrm{d}x=\int_{\mathbb{R}^{d}}e^{-iy\frac{\xi}{\sqrt{n}\bar{\sigma}}}f_{\sigma_{1}}*\cdots*f_{\sigma_{n}}(y)\,\mathrm{d}x\\ &=\mathcal{F}(f_{\sigma_{1}}*\cdots f_{\sigma_{n}})\Big{(}\frac{\xi}{\sqrt{n}\bar{\sigma}}\Big{)}=\prod_{i=1}^{n}\hat{f}_{\sigma_{i}}\Big{(}\frac{\xi}{\sqrt{n}\bar{\sigma}}\Big{)}=\prod_{i=1}^{n}\int_{\mathbb{R}^{d}}e^{-ix\frac{\xi}{\sqrt{n}\bar{\sigma}}}\sigma_{i}^{-d}f\Big{(}\frac{x}{\sigma_{i}}\Big{)}\,\mathrm{d}x\\ &=\prod_{i=1}^{n}\int_{\mathbb{R}^{d}}e^{-iy\frac{\xi\sigma_{i}}{\sqrt{n}\bar{\sigma}}}f(y)\,\mathrm{d}y=\prod_{i=1}^{n}\hat{f}\Big{(}\frac{\xi\sigma_{i}}{\sqrt{n}\bar{\sigma}}\Big{)},\end{split}$
and
$\hat{G}(\xi)=\prod_{i=1}^{n}\hat{G}\Big{(}\frac{\xi\sigma_{i}}{\sqrt{n}\bar{\sigma}}\Big{)}.$
@@
$\left\|\prod_{i=1}^{n}\hat{f}\Big{(}\frac{\xi\sigma_{i}}{\sqrt{n}\bar{\sigma}}\Big{)}\right\|_{L^{1}}\leq\prod_{i=1}^{n}\left\|\hat{f}\Big{(}\frac{\xi\sigma_{i}}{\sqrt{n}\bar{\sigma}}\Big{)}\right\|_{L^{n}}\leq\prod_{i}^{n}\Big{(}\frac{\bar{\sigma}\sqrt{n}}{\sigma^{i}}\Big{)}^{d/n}\left\|\hat{f}\right\|_{L^{n}}\leq n^{d/2}\Big{(}\frac{L}{l}\Big{)}^{d}\left\|\hat{f}\right\|_{L^{n}}^{n},$
@@
$\left|\mathfrak{f}_{n}(x)-G(x)\right|=(2\pi)^{d}\left|\int_{\mathbb{R}^{d}}(\hat{\mathfrak{f}}_{n}(\xi)-\hat{G}(\xi))e^{i\xi x}\,\mathrm{d}\xi\right|\leq(2\pi)^{d}\int_{\mathbb{R}^{d}}\left|\hat{\mathfrak{f}}_{n}-\hat{G}\right|\,\mathrm{d}\xi$

### v3

@@ 3.1 Harris’s Theorem
where $L^{*}$ is the (formal) dual operator of $L$ acting on “nice enough” functions (see for example Hairer (2010), where inequality (14) must be understood in a submartingale sense). We show condition (14), in a suitable sense, for the generator associated to equation (NLFP) in Theorems 3.2 and 3.4. From this we will deduce Hypothesis 1.
@@
A stochastic operator $S$ on $\mathcal{M}$ satisfies a Harris condition on a set $\mathcal{S}\subseteq\Omega$ if there exist $0<\alpha<1$ and $\rho\in\mathcal{P}$ such that
@@
[…]
@@ 3.2 Lyapunov condition
To lighten the notation in the following, we will use $C_{1},C_{2}$ as general constants, whose value may change from line to line. Recall from (1), that for all $i=1,\dots,d$
$\int J(z)z_{i}\,\mathrm{d}z=0.$
@@
where $R_{\alpha}(\xi)$ is the Lagrange reminder for some $\xi(x,z)\in$ $x,x+\varepsilon z]$, the segment that joins $x$ and $x+\varepsilon z$.
@@
$\partial_{i}^{2}\left\langle x\right\rangle^{k}=k\big(\left\langle x\right\rangle^{k-2}+(k-2)x_{i}^{2}\left\langle x\right\rangle^{k-4}\big).$ (15)
For $k\in(0,2)$ we have $\partial_{i}^{2}\varphi(\xi)\leq k$, for every $i,\xi$ . Hence
$\frac{1}{\varepsilon^{2}}\int_{\mathbb{R}^{d}}J_{\varepsilon}(y)(\left\langle x+y\right\rangle^{k}-\left\langle x\right\rangle^{k})\,\mathrm{d}y\leq\frac{k}{2}\int_{\mathbb{R}^{d}}J(z)|z|^{2}\,\mathrm{d}z=k.$
@@
$\begin{split}\partial_{i}^{2}\varphi(\xi)&\leq k(k-1)\langle\xi\rangle^{k-2}\leq k(k-1)\left\langle|x|+\varepsilon|z|\right\rangle_{k-2}\leq C_{1}(\left\langle x\right\rangle^{k-2}+\left\langle z\right\rangle^{k-2}).\end{split}$
@@
$\begin{split}\frac{1}{\varepsilon^{2}}&\int_{\mathbb{R}^{d}}J_{\varepsilon}(y)(\left\langle x+y\right\rangle^{k}-\left\langle x\right\rangle^{k})\,\mathrm{d}y\leq\frac{1}{2}\int_{\mathbb{R}^{d}}J(z)\sum_{i=1}^{d}z_{i}^{2}C_{1}(\langle x\rangle^{k-2}+\langle z\rangle^{k-2})\,\mathrm{d}z\\ &\leq C_{1}\langle x\rangle^{k-2}\frac{1}{2}\int_{\mathbb{R}^{d}}J(z)|z|^{2}\,\mathrm{d}z+C_{1}\int_{\mathbb{R}^{d}}J(z)\langle z\rangle^{k}\,\mathrm{d}z\leq C_{2}\langle x\rangle^{k-2}.\end{split}$
@@
We point out that as $k$ approaches 0, the right hand side of (16) goes to 0 as well, since we don’t have Lyapunov confinement in $L^{1}$ without weight.
@@
To lighten the notation in the following, we will use $C_{1},C_{2}$ as general constants, whose value may change from line to line. As before, for all $i=1,\dots,d$
$\int_{\mathbb{R}^{d}}J(z)z_{i}\,\mathrm{d}z=0.$
As before, for every $\varphi\in\mathcal{C}^{\infty}$, we Taylor expand around $x$, obtaining
@@
where the Lagrange reminder is evaluated at some $\xi(x,z)\in$ $x,x+\varepsilon z]$.
@@
$\partial_{i}^{2}\varphi(\xi)=\frac{ae^{a\left\langle\xi\right\rangle}}{\left\langle\xi\right\rangle^{2}}\Big(a\xi_{i}^{2}+\left\langle\xi\right\rangle-\frac{\xi_{i}^{2}}{\left\langle\xi\right\rangle}\Big)\leq C_{1}e^{a\left\langle\xi\right\rangle}\leq C_{1}e^{a(\left\langle x\right\rangle+\varepsilon\left\langle z\right\rangle)}$
@@
We thus have proved a Lyapunov condition on the generator for $V=\left\langle\cdot\right\rangle^{k}$ and $V=e^{a\left\langle\cdot\right\rangle}$. For these weight functions (and in general for $\mathcal{C}^{2}$ functions), we see that actually Assumption (14) implies Assumption (SLC), which is the statement of the next Theorem.
@@
Let $u_{0}$ be in $\mathcal{D}(L_{\varepsilon})$ by standard semigroup theory $u(t)\in\mathcal{C}^{1}([0,\infty);L^{1}_{k})$, and consequently we compute
$\begin{split}\frac{\mathrm{d}}{\mathrm{d}t}\int_{\mathbb{R}^{d}}u(t,\cdot)\varphi(x)\,\mathrm{d}x&=\int_{\mathbb{R}^{d}}\frac{\mathrm{d}}{\mathrm{d}t}u(t,\cdot)\varphi(x)\,\mathrm{d}x\\ &=\int_{\mathbb{R}^{d}}(Lu)\varphi(x)\stackrel{{\scriptstyle\mathclap{\mbox{*}}}}{{=}}\int_{\mathbb{R}^{d}}u(L^{*}\varphi)\leq\int_{\mathbb{R}^{d}}u(-\lambda\varphi+C)\end{split}$
[…]
@@
$f^{\prime}(t)\leq-\lambda f(t)+C\implies f(t)\leq e^{-\lambda t}f(0)+\frac{C}{\lambda}(1-e^{-\lambda t}).$
Since the $\mathcal{D}(L_{\varepsilon})$ is dense in $L^{1}_{k}$, the claim is true also for any $u_{0}\in L^{1}_{k}$ by classic density argument. ∎
@@ 3.3 An $L^{\infty}$ version of Berry-Esseen’s central limit theorem
Berry-Esseen-type theorems are quantitative versions of the central limit theorem which provide an estimate on the error between the normal distribution and the scaled law. A result of this kind will be crucial for our proof in Section 3.4 of positivity for the nonlocal Fokker-Planck equation. The first version of this theorem was proved independently by Berry (1941) and Esseen (1942) for real, centered i.i.d. random variables $X_{1},\dots X_{n}$ with ${\textrm{Var}}\,(X_{1})=\sigma^{2}$ and $\mathbb{E}(|X_{1}|^{3})=\rho$, obtaining a result of the type
$\sup_{x\in\mathbb{R}}|F_{n}(x)-\Phi(x)|\leq\frac{C\rho}{\sqrt{n}\sigma^{3}}$
where $\Phi$ is the cumulative distribution function (c.d.f.) of the normal distribution, $F_{n}$ is the c.d.f. of $\bar{S}_{n}=\frac{1}{\sigma\sqrt{n}}\sum_{i=1}^{n}X_{i}$, and $C$ is a universal constant depending on the dimension $d$. Several generalizations have been developed, proving the same kind of results for not i.i.d. random vectors in $\mathbb{R}^{d}$, as well as improving the precision of the value of $C$. Recent versions can be found in Bentkus (2005) and Raič (2019).
@@
Then, there exist an integer $N=N(p)$ and a constant $C_{BE}$ depending on $\frac{L}{l},p,d,\rho_{2+s},\left\|f\right\|_{L^{p}}$ such that $\mathfrak{f}_{n}\in L^{\infty}$ for all $n\geq N$ and
@@
where $G$ is the density of the standard normalized Gaussian.
In order to prove this we first present some preliminary results. The first one is a generalization of the identity $a^{n}-b^{n}=\sum_{i=1}^{n}(a-b)a^{i-1}b^{n-i}$, that can be easily proven by induction:
@@
$\max_{|\alpha|=2}|D^{\alpha}\hat{f}(x)-D^{\alpha}\hat{f}(y)|\leq 4\rho_{2+s}|x-y|^{s}.$
@@
The $\mathcal{C}^{2}$-regularity is immediate from Fourier transform properties. We are left to show the Hölder-continuity of the derivatives of order $2$, namely, for every multi-index $\alpha$ such that $|\alpha|=2$,
$\begin{split}&\left|D^{\alpha}\hat{f}(\xi)-D^{\alpha}\hat{f}(\eta)\right|=\left|-\widehat{x^{\alpha}f}(\xi)+\widehat{x^{\alpha}f}(\eta)\right|=\left|\int_{\mathbb{R}^{d}}\Big(e^{-ix\eta}-e^{-ix\xi})x^{\alpha}f(x)\,\mathrm{d}x\right|\\ &\leq\int_{|x|\leq R}|e^{-ix\xi}-e^{-ix\eta}||x|^{2}|f(x)\,\mathrm{d}x+\int_{|x|>R}\left|e^{-ix\xi}-e^{-ix\eta}\right||x|^{2}|f(x)\,\mathrm{d}x\\ &\leq 2|\xi-\eta|\int_{|x|\leq R}|x|^{2+s}|x|^{1-s}f(x)\,\mathrm{d}x+2R^{-s}\int_{|x|>R}|x|^{2+s}f(x)\,\mathrm{d}x\\ &\leq 2\rho_{2+s}(|\xi-\eta|R^{1-s}+R^{-s}).\end{split}$
@@
The last ingredient we need is the following technical Lemma (that is basically (Hauray and Mischler, 2014, lemma 4.8) with a weaker assumption on moments. We denote by $B_{\delta}(x_{0})$, the ball in $\mathbb{R}^{d}$ centered at $x_{0}$ with radius $\delta$. When the ball is centered at 0, we use $B_{\delta}$ as a shorthand.
@@
    $\forall\xi\in B_{\delta}\quad|\hat{f}(\xi)|\leq e^{-\frac{|\xi|^{2}}{4}}$
@@
    Assume additionally that $f\in L^{p}$, with $p\in(1,\infty]$. For any $\delta>0$ there exists $\kappa=\kappa(\rho_{2+s},\delta,\left\|f\right\|_{L^{p}})\in(0,1)$ such that sup_|ξ|δ|^f(ξ)|κ(δ).
@@
    $\begin{split}&\left|\int_{0}^{|\xi|}\Big(\frac{\xi}{|\xi|}\Big)^{T}\Big(H_{\hat{f}}\Big(z\frac{\xi}{|\xi|}\Big)-H_{\hat{f}}(0)\Big)\Big(\frac{\xi}{|\xi|}\Big)(z-|\xi|)\,\mathrm{d}z\right|\\ &\leq\int_{0}^{|\xi|}\left|\!\left|\!\left|(H_{\hat{f}}\Big(z\frac{\xi}{|\xi|}\Big)-H_{\hat{f}}(0)\right|\!\right|\!\right||z-|\xi||\,\mathrm{d}z\\ &\leq 4\rho_{2+s}\int_{0}^{|\xi|}|z|^{s}|z-|\xi||\,\mathrm{d}z=\frac{4\rho_{2+s}|\xi|^{2+s}}{(s+1)(s+2)},\end{split}$
    where $H_{f}(x)$ is the Hessian matrix of $f$ applied at point $x$ and $\left|\!\left|\!\left|\cdot\right|\!\right|\!\right|$ is the spectral norm which is always less or equal than the $\max$- norm used in Lemma 3.8.
    Thus, for small $|\xi|<\delta$ which explicitly depends only on $s$ and $\rho_{2+s}$, we have |^f(ξ)|≤|1-38|ξ|^2|≤e^-|ξ|24.
@@
We can now present the proof of our version of Berry-Esseen theorem, that follows the ideas presented in (Goudon et al., 2002, Thm. 1) and (Hauray and Mischler, 2014, Thm. 4.6).
@@
(NOT THE UPDATED VERSION: $l/L$) First, we notice that $\mathfrak{f}_{n}$ satisfies the hypotheses of Lemma 3.9 for every $n$, and, in particular
$\int_{\mathbb{R}^{d}}\mathfrak{f}_{n}(x)x_{i}x_{k}\,\mathrm{d}x=\delta_{ij}.$
@@
$\begin{split}{\hat{\mathfrak{f}}_{n}}(\xi)&=\int_{\mathbb{R}^{d}}e^{-ix\xi}(\bar{\sigma}^{2}n)^{d/2}f^{\sigma_{1}\dots,\sigma_{n}}(\sqrt{n}\bar{\sigma}x)\,\mathrm{d}x=\int_{\mathbb{R}^{d}}e^{-iy\frac{\xi}{\sqrt{n}\bar{\sigma}}}f_{\sigma_{1}}*\cdots*f_{\sigma_{n}}(y)\,\mathrm{d}x\\ &=\mathcal{F}(f_{\sigma_{1}}*\cdots*f_{\sigma_{n}})\Big(\frac{\xi}{\sqrt{n}\bar{\sigma}}\Big)=\prod_{j=1}^{n}\hat{f}_{\sigma_{j}}\Big(\frac{\xi}{\sqrt{n}\bar{\sigma}}\Big)=\prod_{j=1}^{n}\int_{\mathbb{R}^{d}}e^{-ix\frac{\xi}{\sqrt{n}\bar{\sigma}}}\sigma_{j}^{-d}f\Big(\frac{x}{\sigma_{j}}\Big)\,\mathrm{d}x\\ &=\prod_{j=1}^{n}\int_{\mathbb{R}^{d}}e^{-iy\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}}f(y)\,\mathrm{d}y=\prod_{j=1}^{n}\hat{f}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big),\end{split}$
and, since $G$ is the standard gaussian, its Fourier transform satisfies
$\hat{G}(\xi)=\prod_{j=1}^{n}\hat{G}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big).$
@@
$\left\|\prod_{j=1}^{n}\hat{f}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)\right\|_{L^{1}}\leq\prod_{j=1}^{n}\left\|\hat{f}\Big(\frac{\xi\sigma_{j}}{\sqrt{n}\bar{\sigma}}\Big)\right\|_{L^{n}}=\prod_{j=1}^{n}\Big(\frac{\bar{\sigma}\sqrt{n}}{\sigma_{j}}\Big)^{d/n}\left\|\hat{f}\right\|_{L^{n}}\leq n^{d/2}\Big(\frac{L}{l}\Big)^{d}\left\|\hat{f}\right\|_{L^{n}}^{n},$
@@
$\left|\mathfrak{f}_{n}(x)-G(x)\right|=(2\pi)^{-d}\left|\int_{\mathbb{R}^{d}}(\hat{\mathfrak{f}}_{n}(\xi)-\hat{G}(\xi))e^{i\xi x}\,\mathrm{d}\xi\right|\leq(2\pi)^{-d}\int_{\mathbb{R}^{d}}\left|\hat{\mathfrak{f}}_{n}-\hat{G}\right|\,\mathrm{d}\xi$